OpenStudy (anonymous):

Need help with 1,2,13,14

OpenStudy (anonymous):

mmmmm what with does #

OpenStudy (anonymous):

for the first one is x in the corner of the cone?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

or is it the question mark?

OpenStudy (anonymous):

that stands for x

OpenStudy (anonymous):

ok so for the first one ibelieve is B.. dont trust me

OpenStudy (anonymous):

why b? can u explain

OpenStudy (anonymous):

as you seee there is 28+8=36 and x+ 7=x

OpenStudy (anonymous):

soo the line were x is bigger than the other one so the answer has to be greater

OpenStudy (anonymous):

so 32+7=39 and 34.1+7= 41.1

OpenStudy (anonymous):

so it could be A or B but... is you see the line in the bottom there is 8 and 7 and they look bigg

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

and they are not equal or the same if not they would have the little lines in the line so most likely, is either A or B but i go more with B

OpenStudy (anonymous):

i ahve to go no battery sorry

OpenStudy (anonymous):